# Adaptive Exploration for Multi-Reward Multi-Policy Evaluation

### v2

@@ 4.2 Sample Complexity Lower Bound
As a consequence of Proposition 4.1, the analysis of the sample complexity must necessarily take into account the set of rewards. We let ${\cal R}_{\pi}^{\epsilon}=\{r\in{\cal R}_{\pi}:{\rm Alt}_{\pi,r}^{\epsilon}(M)\neq\emptyset\}$ be the set of rewards for which the corresponding set of confusing models is non-empty.
To derive the sample complexity lower bound, we also define the characteristic time $T_{\epsilon}(\omega;M)$ of a stationary state-action distribution $\omega$ under $M$:
$T_{\epsilon}(\omega;M)^{-1}\coloneqq\inf_{\pi\in\Pi,r\in{\cal R}_{\pi}^{\epsilon},M_{r}^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)}\mathbb{E}_{\omega}[{\rm KL}_{P|P_{r}^{\prime}}(s,a)],$ (2)
In this optimization problem, we seek an alternative model $M^{\prime}$ that is confusing, i.e., that minimizes the statistical difference from the true process $M$ under some pair $(\pi,r)$. Conversely, the sampling distribution $\omega$ serves to gather evidence distinguishing $M$ from any such alternative. Hence, the reciprocal $T_{\epsilon}(\omega;M)^{-1}$ can be interpreted as the information‐gathering rate under $\omega$ in $M$.
@@
$T_{\epsilon}^{\star}(M)=\min_{\omega\in\Omega(M)}T_{\epsilon}(\omega;M),$ (3)
where we denote by $\omega_{\rm opt}$ the solution to the optimization in Eq. (3), and $\Omega(M)$ is defined as the set of state-action distributions satisfying the forward equations: $\Omega(M)\coloneqq\Big{\{}\omega\in\Delta({\cal S}\times{\cal A}):\sum_{a}\omega(s,a)=\sum_{s^{\prime},a^{\prime}}P(s|s^{\prime},a^{\prime}),\omega(s^{\prime},a^{\prime})\;\forall s\Big{\}}.$ We have the following asymptotic lower bound, and the proof of the theorem can be found in Section B.2.1.
@@
For any $(\epsilon,\delta)$-PAC algorithm we have $\liminf_{\delta\to 0}\frac{\mathbb{E}[\tau]}{\log(1/\delta)}\geq T_{\epsilon}^{\star}(M).$ (4)
As discussed earlier, the optimal exploration strategy $\omega_{\rm opt}$, in terms of state-action visitations, maximizes information gain per time-step required to distinguish between the true model $M$ and a confusing one.
@@
However, it it may seem counterintuitive to use an exploration strategy that depends on the true model itself $M$, since it is unknown. In practice, we use the empirical estimate of the MDP $M_{t}$ at time-step $t$ to derive an exploration strategy. We defer this discussion to Section 5.
@@
Lemma 4.4 (Impossibility).
Let $T^{\pi_{e}}$ be the set of transient states (i.e., not recurrent) under a behavior policy $\pi_{e}$. Assume that $T^{\pi_{e}}\neq\emptyset$: if there exists $s_{c}\in T^{\pi_{e}},\pi\in\Pi,r\in{\cal R}_{\pi}$ such that $\|\rho_{r}^{\pi}(s_{c})\|_{\infty}>2\epsilon/\gamma$, then $\pi_{e}$ is not optimal, in the sense that $T_{\epsilon}(d^{\pi_{e}};M)^{-1}=0,$ (5) for any stationary distribution $d^{\pi_{e}}$ (of the chain $P^{\pi_{e}}$) induced by $\pi_{e}$.
The proof for this lemma can be found in Section B.2.2. This result is general, and also applies to to single-policy single-reward cases. Moreover, we understand that an optimal exploration policy needs to visit states with large deviation gaps $\|\rho_{r}^{\pi}(s)\|_{\infty}$ infinitely often. As we see in the next section, we can devise an exploration policy based on this intuition.
@@ 4.2.1 Convexity of the Lower Bound
We find that it is hard to directly optimize $T_{\epsilon}(\omega;M)$, since the optimization the set of alternative models may be non-convex. Observe that the set ${\rm Alt}_{\pi,r}^{\epsilon}(M)$ can be seen as the union of two sets $\{M^{\prime}:\max_{s}V_{M_{r}}^{\pi}(s)-V_{M_{r}^{\prime}}^{\pi}(s)\geq 2\epsilon\}$ and $\{M^{\prime}:\max_{s}V_{M_{r}^{\prime}}^{\pi}(s)-V_{M_{r}}^{\pi}(s)\geq 2\epsilon\}$. The convexity of these sets depend on $(\pi,r,P)$ and, even if convex, may be disjoint. The following example illustrates this aspect.
@@ 4.3 Relaxed Characteristic Time
We proceed with finding a convex relaxation of $T_{\epsilon}(\omega;M)$ that holds for all distributions $\omega\in\Delta({\cal S}\times{\cal A})$. Such relaxation not only upper bounds $T_{\epsilon}(\omega;M)$ in terms of ${\cal R}_{\pi}$ (instead of ${\cal R}_{\pi}^{\epsilon}$), but also allows to better understand the scaling of $T_{\epsilon}^{\star}$. The proof can be found in Section B.2.3.
@@
For all $\omega\in\Delta({\cal S}\times{\cal A})$ we have $T_{\epsilon}(\omega;M)\leq U_{\epsilon}(\omega;M)$, where $U_{\epsilon}(\omega;M)\coloneqq\sup_{\pi\in\Pi,r\in{\cal R}_{\pi}}\max_{s\in{\cal S}}\frac{\gamma^{2}\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}}{4\epsilon^{2}(1-\gamma)^{2}\omega(s,\pi(s))},$ (6) is convex in $\omega$. Let $U_{\epsilon}^{\star}(M)=U_{\epsilon}(\omega^{\star};M)$ be the optimal rate, where $\omega^{\star}=\operatorname*{arg\,min}_{\omega\in\Omega(M)}U_{\epsilon}(\omega;M)$ is the $U_{\epsilon}$-optimal visitation strategy.
@@
To better understand the scaling, for example, in the generative setting, choosing a uniform distribution $\omega(s,\pi(s))=1/S$ yields a scaling of $O\left(\max_{\pi,r}\frac{\gamma^{2}\,\lvert\mathcal{S}\rvert\,\max_{s}\|\rho_{r}^{\pi}(s)\|^{2}}{4\epsilon^{2}(1-\gamma)^{2}}\right)$. Since $\|\rho_{r}^{\pi}(s)\|_{\infty}\leq 1/(1-\gamma)$, we obtain a minimax scaling of $O\left(\frac{\gamma^{2}|\mathcal{S}|}{4\epsilon^{2}(1-\gamma)^{4}}\right),$ which is independent of the number of policies or rewards. Note that this scaling can be misleading: because the sample complexity scales according to $\frac{\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}}{\omega(s,\pi(s))}$, a small $\omega$ does not necessarily increase sample complexity if $\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}$ is also small. In Section B.2.5 we depict the scaling on a simple example for two types of reward sets: reward-free (i.e., the entire set of rewards) and the circle $\|r\|_{2}\leq 1$.
@@ 4.3.1 Optimization over Convex Sets of Rewards and Reward-Free Policy Evaluation
For a convex reward set $\mathcal{R}_{\pi}$, the maximization over the rewards in Theorem 4.6 can be cast as a mixed-integer linear program (MILP). In practice, as shown in Section B.2.5, the same optimum is obtained more efficiently by solving a finite collection of ordinary convex programs. Therefore, the optimization problem can be formulated within the framework of disciplined convex programming (DCP) (Grant et al., 2006).
@@
$K^{\pi}(s)=I-\mathbf{1}P(s,\pi(s))^{\top},\quad G^{\pi}=(I-\gamma P^{\pi})^{-1},$
and $\Gamma_{i,j}^{\pi}(s)$ can be interpreted as the expected discounted number of visits to state $j$ starting from $i$ after subtracting the expected number of visits to $j$ starting from $s^{\prime}\sim P(s,\pi(s))$. In other words, $\Gamma^{\pi}(s)$ is the analogue of the deviation $\rho$ in terms of discounted number of visits.
@@
If ${\cal R}_{\pi}=[0,1]^{S}$, we have that $U_{\epsilon}(\omega;M)=\max_{\pi,s,s^{\prime}}\frac{\gamma^{2}\left[\max\left(\Gamma_{+}^{\pi}(s,s^{\prime}),\Gamma_{-}^{\pi}(s,s^{\prime})\right)\right]^{2}}{4\epsilon^{2}(1-\gamma)^{2}\omega(s,\pi(s))}.$
@@ 5 MR-NaS for Policy Evaluation
In this section show how to adapt the MR-NaS (Multi-Reward Navigate and Stop) algorithm (Russo & Vannella, 2024) for multi-reward multi-policy evaluation based on the results from the previous section. MR-NaS (Algorithm 1) is designed with 2 key components: (1) a sampling rule and (2) a stopping rule. We now discuss each of these.
@@
The key idea is to sample according to the policy induced by $\omega^{\star}=\operatorname*{arg\,min}_{\omega\in\Omega(M)}U_{\epsilon/2}(\omega;M)$. Indeed, sampling actions according to $\pi^{\star}(a|s)=\omega^{\star}(s,a)/\sum_{b}\omega^{\star}(s,b)$ guarantees optimality with respect to $U_{\epsilon/2}^{\star}$, as the solution $\omega^{\star}$ matches the rate in Theorem 4.3. The factor $\epsilon/2$ arises from the lower bound analysis that requires $2\epsilon$-separation. By tightening the accuracy to $\epsilon/2$, we ensure the $(\epsilon,\delta)$-PAC guarantee, which would otherwise be hard to prove. This results in an additional constant factor $4$ in the sample complexity.
However, $\omega^{\star}$ cannot be computed without knowledge of the MDP $M$. As in previous works (Garivier & Kaufmann, 2016; Al Marjani et al., 2021), we employ the certainty equivalence principle (CEP): plug-in the current estimate at time $t$ of the transition function and compute $\omega_{t}^{\star}$. The allocation $\omega_{t}^{\star}$ rapidly eliminates models confusing for $M_{t}$, efficiently determining whether the true model $M$ is non-confusing–motivating our use of the CEP.
@@
2:     Compute $\omega_{t}^{\star}=\operatorname*{arg\,inf}_{\omega\in\Omega(M_{t})}U_{\epsilon/2}(\omega;M_{t})$.
@@
In summary, the algorithm proceeds as follows: at each time-step $t$ the agent computes the optimal visitation distribution $\omega_{t}^{\star}=\operatorname*{arg\,inf}_{\omega\in\Omega(M_{t})}U_{\epsilon/2}(\omega;M_{t})$ with respect to $M_{t}$, the estimate of the MDP (which is, practically speaking, the estimate of the transition function). The policy $\pi_{t}^{\star}(a|s)$ induced by $\omega_{t}^{\star}$ is mixed with a forcing policy $\pi_{f,t}(\cdot|s)$ (e.g., a uniform distribution over actions or a distribution that encourages to select under-sampled actions; see also (Russo & Vannella, 2024) or Section C.3 for more details) that guarantees all actions are sampled infinitely often. The mixing factor $\varepsilon_{t}$ can be chosen as $\varepsilon_{t}=1/N_{t}(s_{t})$, where $N_{t}(s)$ is the number of visits of state $s$ up to time $t$. The resulting exploration policy, $\pi_{t}$, is used to sample an action $a_{t}$. Upon observing the next state, the transition function is updated using the empirical average.
Remark 5.1.
At first glance, the optimization problem $\min_{\omega\in\Omega(M_{t})}U_{\epsilon/2}(\omega;M)$ might appear to overlook actions $a\neq\pi(s)$. However, these actions are accounted for through the forward constraint $\omega\in\Omega(M_{t})$.
@@
Lastly, the method stops whenever sufficient evidence has been gathered to obtain $(\epsilon,\delta)$-PAC guarantees. These requires approximately $U_{\epsilon/2}^{\star}(M)\log(1/\delta)$ samples (by inspecting Theorems 4.3 and 4.6).

### v3

@@ 4.2 Sample Complexity Lower Bound
As a consequence of Proposition 4.1, the analysis of the sample complexity must necessarily take into account the reward sets. However, for clarity, we adopt the assumption that for every state there exists a confusing model (see Prop. 4.1). While not strictly necessary (one could work with ${\cal R}_{\pi}^{\epsilon}$ instead), this assumption simplifies our analysis.
Assumption 4.3.
For every state $s\in{\cal S}$ there exists $\pi\in\Pi,r\in{\cal R}_{\pi}$ such that $\|\rho_{r}^{\pi}(s)\|_{\infty}>\frac{2\epsilon}{\gamma}$.
To derive the sample complexity lower bound, we define the characteristic time $T_{\epsilon}(\omega;M)$ of a stationary state-action distribution $\omega$ under $M$:
$T_{\epsilon}(\omega;M)^{-1}\coloneqq\inf_{\pi\in\Pi,r\in{\cal R}_{\pi},M_{r}^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)}\mathbb{E}_{\omega}[{\rm KL}_{P|P_{r}^{\prime}}(s,a)],$ (2)
where $(s,a)\sim\omega$. In this optimization problem, we seek an alternative model $M^{\prime}$ that is confusing, i.e., that minimizes the statistical difference from the true process $M$ under some pair $(\pi,r)$. Conversely, the sampling distribution $\omega$ serves to gather evidence distinguishing $M$ from any such alternative. Hence, the reciprocal $T_{\epsilon}(\omega;M)^{-1}$ can be interpreted as the information‐gathering rate under $\omega$ in $M$.
@@
$T_{\epsilon}^{\star}(M)=\inf_{\omega\in\Omega(M)}T_{\epsilon}(\omega;M),$ (3)
where we denote by $\omega_{\rm opt}\in\overline{\Omega}(M)$ a solution to the optimization in Eq. (3), and $\Omega(M)$ is defined as the following set of stationary state-action distributions: $\Omega(M)\coloneqq\Big{\{}\omega\in\Delta({\cal S}\times{\cal A}):(\sum_{a}\omega(s,a)=\sum_{s^{\prime},a^{\prime}}P(s|s^{\prime},a^{\prime})\omega(s^{\prime},a^{\prime})\,\forall s)\hbox{ and }\omega(s,a)>0\;\forall(s,a)\big{\}}$. We have the following asymptotic lower bound, and the proof of the theorem can be found in Section B.2.1.
@@
Under 4.3, for any $(\epsilon,\delta)$-PAC algorithm we have $\liminf_{\delta\to 0}\frac{\mathbb{E}[\tau]}{\log(1/\delta)}\geq T_{\epsilon}^{\star}(M).$ (4)
As discussed earlier, an exploration strategy matching the information gain $T_{\epsilon}^{\star}(M)$ maximizes the evidence collected per time-step required to distinguish between the true model $M$ and a confusing one.
@@
However, it may seem counterintuitive to use an exploration strategy that depends on the true model itself $M$, since it is unknown. In practice, we use the empirical estimate of the MDP $M_{t}$ at time-step $t$ to derive an exploration strategy. We defer this discussion to Section 5.
@@
Lemma 4.5.
Let $T^{\pi_{e}}$ be the set of transient states (i.e., not recurrent) under a behavior policy $\pi_{e}$. Assume that $T^{\pi_{e}}\neq\emptyset$. If there exists $s_{c}\in T^{\pi_{e}},\pi\in\Pi,r\in{\cal R}_{\pi}$ such that $\|\rho_{r}^{\pi}(s_{c})\|_{\infty}>2\epsilon/\gamma$, then $\pi_{e}$ is not optimal, in the sense that $T_{\epsilon}(d^{\pi_{e}};M)^{-1}=0,$ (5) for any stationary distribution $d^{\pi_{e}}$ (of the chain $P^{\pi_{e}}$) induced by $\pi_{e}$.
The proof for this lemma can be found in Section B.2.2. This result is general, and also applies to single-policy single-reward cases. Moreover, we understand that an optimal exploration policy needs to frequently sample states with large deviation gaps $\|\rho_{r}^{\pi}(s)\|_{\infty}$. This intuition will be used to devise an exploration policy.
@@ 4.2.1 Convexity of the Lower Bound
We find that it is hard to directly optimize $T_{\epsilon}(\omega;M)$, since the optimization over the set of alternative models may be non-convex. Observe that the set ${\rm Alt}_{\pi,r}^{\epsilon}(M)$ can be seen as the union of two sets $\{M^{\prime}:\max_{s}V_{M_{r}}^{\pi}(s)-V_{M_{r}^{\prime}}^{\pi}(s)\geq 2\epsilon\}$ and $\{M^{\prime}:\max_{s}V_{M_{r}^{\prime}}^{\pi}(s)-V_{M_{r}}^{\pi}(s)\geq 2\epsilon\}$. The convexity of these sets depends on $(\pi,r,P)$ and, even if convex, may be disjoint. The following example illustrates this aspect.
@@ 4.3 Relaxed Characteristic Time
We proceed with finding a convex relaxation of $T_{\epsilon}(\omega;M)$ that holds for all distributions $\omega\in\Delta({\cal S}\times{\cal A})$. Such relaxation not only upper bounds $T_{\epsilon}(\omega;M)$ in terms of ${\cal R}_{\pi}$ (instead of ${\cal R}_{\pi}^{\epsilon}$), but also allows us to better understand the scaling of $T_{\epsilon}^{\star}$. The proof can be found in Section B.2.3.
@@
For all $\omega\in\Delta({\cal S}\times{\cal A})$ we have $T_{\epsilon}(\omega;M)\leq U_{\epsilon}(\omega;M)$, where $U_{\epsilon}(\omega;M)\coloneqq\sup_{\pi\in\Pi,r\in{\cal R}_{\pi}}\max_{s\in{\cal S}}\frac{\gamma^{2}\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}}{2\epsilon^{2}(1-\gamma)^{2}\omega(s,\pi(s))},$ (6) is convex in $\omega$. Let $U_{\epsilon}^{\star}(M)=U_{\epsilon}(\omega^{\star};M)$ be the optimal rate, where $\omega^{\star}\in\operatorname*{arg\,inf}_{\omega\in\overline{\Omega(M)}}U_{\epsilon}(\omega;M)$ is an $U_{\epsilon}$-optimal allocation.
@@
To better understand the scaling, for example, in the generative setting, choosing a uniform distribution $\omega(s,\pi(s))=1/S$ yields a scaling of $O\left(\max_{\pi,r}\frac{\gamma^{2}\,\lvert\mathcal{S}\rvert\,\max_{s}\|\rho_{r}^{\pi}(s)\|^{2}}{2\epsilon^{2}(1-\gamma)^{2}}\right)$. Since $\|\rho_{r}^{\pi}(s)\|_{\infty}\leq 1/(1-\gamma)$, we obtain a minimax scaling of $O\left(\frac{\gamma^{2}|\mathcal{S}|}{2\epsilon^{2}(1-\gamma)^{4}}\right),$ which is independent of the number of policies or rewards. Note that this scaling can be misleading: because the sample complexity scales according to $\frac{\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}}{\omega(s,\pi(s))}$, a small $\omega$ does not necessarily increase sample complexity if $\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}$ is also small. In Section B.2.5 we depict the scaling on a simple example for two types of reward sets: reward-free (i.e., the entire set of rewards) and the circle $\|r\|_{2}\leq 1$.
@@ 4.3.1 Optimization over Convex Sets of Rewards and Reward-Free Policy Evaluation
For a convex reward set $\mathcal{R}_{\pi}$, the maximization over the rewards in Theorem 4.7 can be cast as a mixed-integer linear program (MILP). In practice, as shown in Section B.2.5, the same optimum is obtained more efficiently by solving a finite collection of ordinary convex programs.
@@
$K^{\pi}(s)=I-\mathbf{1}P(s,\pi(s))^{\top},\quad G^{\pi}=(I-\gamma P^{\pi})^{-1}.$
$\Gamma_{i,j}^{\pi}(s)$ can be interpreted as the expected discounted number of visits to state $j$ starting from $i$ after subtracting the expected number of visits to $j$ starting from $s^{\prime}\sim P(s,\pi(s))$. In other words, $\Gamma^{\pi}(s)$ is the analogue of the deviation $\rho$ in terms of discounted number of visits.
@@
If ${\cal R}_{\pi}=[0,1]^{S}\;\forall\pi\in\Pi$, then $U_{\epsilon}(\omega;M)=\max_{\pi,s,s^{\prime}}\frac{\gamma^{2}\left[\max\left(\Gamma_{+}^{\pi}(s,s^{\prime}),\Gamma_{-}^{\pi}(s,s^{\prime})\right)\right]^{2}}{2\epsilon^{2}(1-\gamma)^{2}\omega(s,\pi(s))}.$
@@ 5 MR-NaS for Policy Evaluation
In this section we show how to adapt the MR-NaS (Multi-Reward Navigate and Stop) algorithm (Russo & Vannella, 2024) for multi-reward multi-policy evaluation based on the results from the previous section. MR-NaS (Algorithm 1) is a simple extension of NaS (Al Marjani et al., 2021), and is designed with 2 key components: (1) a sampling rule and (2) a stopping rule. We now discuss each of these.
@@
The key idea is to sample according to the policy induced by $\omega^{\star}\in\operatorname*{arg\,inf}_{\omega\in\overline{\Omega(M)}}U_{\epsilon/2}(\omega;M)$. Indeed, sampling actions according to $\pi^{\star}(a|s)=\omega^{\star}(s,a)/\sum_{b}\omega^{\star}(s,b)$ guarantees optimality with respect to $U_{\epsilon/2}^{\star}$, as the solution $\omega^{\star}$ matches the relaxed rate in Theorem 4.4. The factor $\epsilon/2$ arises from the lower bound analysis that requires $2\epsilon$-separation. By tightening the accuracy to $\epsilon/2$, we ensure the $(\epsilon,\delta)$-PAC guarantee, which would otherwise be hard to prove. This results in an additional constant factor $4$ in the sample complexity.
However, $\omega^{\star}$ cannot be computed without knowledge of the MDP $M$. As in previous works (Garivier & Kaufmann, 2016; Al Marjani et al., 2021), we employ the certainty equivalence principle (CEP): plug in the current estimate at time $t$ of the transition function and compute $\omega_{t}^{\star}$. The allocation $\omega_{t}^{\star}$ rapidly eliminates models confusing for $M_{t}$, efficiently determining whether the true model $M$ is non-confusing–motivating our use of the CEP.
@@
To simplify the analysis, we make the following assumption:
Assumption 5.1.
The solution $\omega^{\star}$ is unique and lies within the open set $\Omega(M)$.
Such assumption prevents: (1) the awkward case where a solution may belong to the boundary of $\overline{\Omega}(M)$; (2) multiple optimal solutions forming a convex set. This latter problem can be addressed using alternative techniques as in (Jedra & Proutiere, 2020; Russo et al., 2025; Russo & Proutiere, 2023b), which use regularization or the idea of tracking a convex combination of past solutions.
In summary, the algorithm proceeds as follows: at each time-step $t$ the agent computes the optimal visitation distribution $\omega_{t}^{\star}=\operatorname*{arg\,inf}_{\omega\in\overline{\Omega(M_{t})}}U_{\epsilon/2}(\omega;M_{t})$ with respect to $M_{t}$, the estimate of the MDP (which is, practically speaking, the estimate of the transition function). The policy $\pi_{t}^{\star}(a|s)$ induced by $\omega_{t}^{\star}$ is mixed with a forcing policy $\pi_{f,t}(\cdot|s)$ (e.g., a uniform distribution over actions or a distribution that encourages to select under-sampled actions; see also (Russo & Vannella, 2024) or Section C.3 for more details) that guarantees all actions are sampled infinitely often. The mixing factor $\varepsilon_{t}$ can be chosen as $\varepsilon_{t}=1/\max(1,N_{t}(s_{t}))$, where $N_{t}(s)$ is the number of visits of state $s$ up to time $t$. The resulting exploration policy, $\pi_{t}$, is used to sample an action $a_{t}$, which, together with the communicating assumption and a forcing policy, yields an ergodic chain. Upon observing the next state, the transition function is updated using the empirical average.
@@
2:  Compute $\omega_{t}^{\star}=\operatorname*{arg\,inf}_{\omega\in\overline{\Omega(M_{t})}}U_{\epsilon/2}(\omega;M_{t})$.
@@
Lastly, the method stops whenever sufficient evidence has been gathered to obtain $(\epsilon,\delta)$-PAC guarantees. This requires approximately $U_{\epsilon/2}^{\star}(M)\log(1/\delta)$ samples (by inspecting Theorems 4.4 and 4.7).